MICROWAVE SPECTROSCOPY OF BUTAN-2-OL - CONFORMATIONS, COMPLEXES AND CLUSTERS.

\underline{ADRIAN K. KING}, BRIAN J. HOWARD, Physical and Theoretical Chemistry Department, University of Oxford, South Parks Road, Oxford, United Kingdom, OX1 3QZ..

The first rotationally-resolved spectrum of a van der Waals complex formed between two chiral molecules will be presented. The complex is a hetero-chiral dimer of butan-2-ol and the spectrum was observed by Fourier Transform Microwave Spectroscopy in a supersonic expansion.^{a} The results of the ab initio calculations used to identify the dimer will also be presented. The microwave spectrum of a complex formed between a butan-2-ol molecule and an argon atom will also be discussed, including a limited amount of structural analysis.^{b}This too will be accompanied by an ab initio study. Finally, mention will be made of the detection of six of the nine possible conformational isomers of the butan-2-ol molecule in expansions of argon and helium.^{c} There are still a large number of unassigned lines in the spectrum of butan-2-ol, some of which must be due to the homo-chiral dimer. Work to assign them is ongoing, any new results will be presented.

aA. K. King and B. J. Howard Chem. Phys. Lett. \underline{348}(343), 2001.

bA. K. King and B. J. Howard J. Mol. Spectrosc., Submitted for publication.

cA. K. King and B. J. Howard J. Mol. Spectrosc. \underline{201/1}(38), 2001.